OpenStudy (anonymous):

MEDAL FOR HELP NOT ANSWERS The perimeter of a rectangle is 126 ft. The ratio of its length to its width is 5 : 4. What are the dimensions of the rectangle? 30 ft by 33 ft 35 ft by 28 ft 6 ft by 60 ft 3 ft by 60 ft

OpenStudy (johnweldon1993):

Well we know that a perimeter is the sum of all the sides so for a rectangle we have \[\large L + L + W + W = 126\] or \[\large 2L + 2W = 126\] now given the ratio...we can say that for every 4 units of length for the width there are...there are 5 units of length for the length of the rectangle...so \[\large 4L = 5W\] or \[\large L = \frac{5}{4}W\] now if we plug that into our equation above we have \[\large 2(\frac{5}{4}W) + 2W = 126\] \[\large \frac{10}{4}W + 2W = 126\] \[\large 10W + 8W = 504\] \[\large 18W = 504\] \[\large W = 28\] and then just solve for L using the fact that W = 28